OpenStudy (anonymous):

solve for the solutionsm set of the following expressions using completing the square: x squared + 9x + 18=0 2x squared + 12x + 4 =0 x squared + 6x-8=0 4 x squared - x - 3=0 what will be the answer .. ?

OpenStudy (campbell_st):

x^2 + 9x + 18 = 0 you need to calculate (9/2)^2 and add to both sides of the equation (x^2 + 9x + 81/4) = 81/4 - 18 or (x + 9/2)^2 = 9/4 take the square root of both sides \[x + \frac{9}{2} = \pm \frac{3}{2}\] all you need to do is subtract 9/2 from both sides for the solutions